ginger made a scale drawing of the brownell carillon tower. the scale was 1:120. if the tower in the drawing is 10.6 inches tall, how tall is the actual tower?
Answer by stanbon(75887)

ginger made a scale drawing of the brownell carillon tower. the scale was 1:120. if the tower in the drawing is 10.6 inches tall, how tall is the actual tower?
--------------------
x/10.6 = 120/1
x = 10.6*120
x = 1272 ft.
=======================
Cheers,
Stan H.
=======================